OpenStudy (anonymous):

simplify each expression 3lnx+2lnx

OpenStudy (anonymous):

52nx i think

OpenStudy (anonymous):

is this supposed to be algebra

OpenStudy (anonymous):

i mean ln(x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can't help

OpenStudy (anonymous):

\[3\ln x + 2\ln x= \ln x ^{3}x ^{2}\]\[\ln x ^{3}x ^{2}=\ln x ^{5}\]